SIMULATION OF FREE→FREE ABSORPTION SPECTRA AND THE CALCULATION OF INTERACTION POTENTIALS FOR ALKALI-RARE GAS ATOM PAIRS

\underline{J. DARBY HEWITT}, THOMAS M. SPINKA, JASON. D. READLE, and J. GARY EDEN, Laboratory for Optical Physics and Engineering University of Illinois at Urbana Champaign Champaign, IL 61820.

We have simulated free→free (X^{2}Σ_{1/2}^{+} →B^{2}Σ_{1/2}^{+}) absorption spectra for alkali-rare gas pairs. By comparing simulation results with experimental data, we have been able to iteratively determine the form for the B^{2}Σ_{1/2}^{+} interaction potential for the system for a range in internuclear separation of 1.5–20 Å. Simulation methods will be presented, as will our results pertaining to Cs-Ar.